If Ernest is suppose to have 75 in sic courses to obtain an A but he has 73 in 5 courses, then what should be his score in the 6 course to still obtain an A?
1
Expert's answer
2020-03-03T10:59:42-0500

Just write the expression that describes his 75% in six courses required to obtain A:


"75=\frac{n_1+n_2+...+n_6}{6},"

where n1, n2, ..., n6 - his individual mark for each course.


75% is the average percentage he wants to obtain for 6 courses whatever the individual mark is. 73% he obtained for 5 courses, and we don't care what each individual mark during 5 courses was:


"73=\frac{n_1+n_2+...+n_5}{5}."

We need to find n6 to tell him what mark he must obtain. From the first expression it is


"n_6=75\cdot6-(n_1+n_2+...+n_5),"


from the second equation we have


"n_1+n_2+...+n_5=73\cdot5."

Substitute this to the third equation:


"n_6=75\cdot6-73\cdot5=85."


He needs to obtain 85% to get A.